OpenStudy (anonymous):

A watch costing $80 is on sale for $56. Find the discount rate.

OpenStudy (anonymous):

find how much percent is 56 of 80

hero (hero):

a - %a = b a = original price b = sales price 80 - %80 = 56 80 - 56 = %80 24 = %80 24/80 = % .3 = % 30 = %

OpenStudy (anonymous):

@Hero answer is in case where it says 30% off. my answer is when it says: " now only 70%"

hero (hero):

The discount rate is definitely not 70%

OpenStudy (anonymous):

:( i'm confused

OpenStudy (anonymous):

it depends on what you understand by discount rate....

hero (hero):

I used percent as a variable. That's probably why you're confused.

hero (hero):

@myko, 70% is not the discount rate.

OpenStudy (anonymous):

80*0.7=56 80-(80*0.3)=56

OpenStudy (anonymous):

i'm confused about all of it. I thought u had to set it up as S=C-rC

hero (hero):

Right, @myko, notice that discount means to "subtract", where in your case, you have "multiplied", rather than subtract.

OpenStudy (anonymous):

maybe...:) I don't go much shopping...:)

hero (hero):

@joyce153, your formula is the same as mine S = b C = a r = %

OpenStudy (anonymous):

Wikipedia says: Rate (mathematics), a specific kind of ratio, in which two measurements are related to each other @Hero

hero (hero):

@myko, why are you telling me things I already know?

hero (hero):

You must never have heard of the term percentage rate.

OpenStudy (anonymous):

sry, didn't whant to barden you...

OpenStudy (anonymous):

?

hero (hero):

Here, we're only looking for the discount rate. I provided the steps necessary to solve the problem above @joyce153

OpenStudy (anonymous):

so i write it like this? 56=80-80r

hero (hero):

Yes, that's how you start off.

hero (hero):

Now, you have to isolate r

OpenStudy (anonymous):

i did 56-80=80-80-80r -24=-80r -24/-80=-80r/-80 & i got 3.3=r

hero (hero):

It would be better if you add 80r to both sides, then subtract 56 from both sides to get 80r = 80 - 56 Then get 80r = 24 Then divide both sides by 80 to get r = 24/80 That reduces to r = 3/10 which is the same thing as 30/100 or 30%

hero (hero):

@szemespeter, we don't use commas as decimals in America.

OpenStudy (phi):

did 56-80=80-80-80r -24=-80r -24/-80=-80r/-80 & i got 3.3=r you were doing ok until the last line -24/-80 = 0.3 so 0.30 = r now change 0.30 to percent by multiplying this fraction by 100 you get 0.30*100= 30%

OpenStudy (anonymous):

@phi how did u get 0.3 from -24/-80? Because when I put that in the calculator it =3.333333333

OpenStudy (anonymous):

if anyone ells can answer that Q let me know.

hero (hero):

You either inputted something wrong or you need to re-calibrate your calculator.

OpenStudy (anonymous):

ok i was doing -80 div -24=3.333333333 -24/-80=3/10

OpenStudy (phi):

type this into the google search window: 24/80= as a sanity check, notice that 80/80 is 1, so 24/80 is smaller than 1 240/80 = 3 how you get 3.3333 is a total mystery.

OpenStudy (phi):

what is div ?

OpenStudy (anonymous):

\[\div\]

OpenStudy (phi):

ohh, you did it upset down!

OpenStudy (anonymous):

ok i think get it now!!

OpenStudy (anonymous):

ok i got it now!! THX!!!

hero (hero):

Wait a second, I help you with the question, but you give phi the medal? :/

OpenStudy (anonymous):

i can't give u both a medal or i would.